OpenStudy (anonymous):

Write an equation for the line in slope-intercept form, given that the line passes through the point (2,3) and (3,4)

OpenStudy (anonymous):

first find the slope, can you do that?

OpenStudy (anonymous):

i don't think i can, im kinda dumb when it comes to math! :(

OpenStudy (anonymous):

now, i know you , you must know somehting about how to get he slope if you are asking these kinds of questions

OpenStudy (anonymous):

the formula for fnding the equation of a line given two points is y2-y1divide by x2-x1 = y-y1divide by x-x1 4-3 divide by 3-2 = y-3 divide by x-2 1=y-3 divide by x-2 x-2 = y-3 x+1 = y